OpenStudy (anonymous):

Find the unknown number. Six added to 3 times a number is equal to 4 less than 5 times the number. 10 5 –5 –10

OpenStudy (anonymous):

-5

OpenStudy (jojo4eva):

the answer is 5

OpenStudy (anonymous):

lol no, 9 times 5 is 45

OpenStudy (anonymous):

Lol uhhh

OpenStudy (jojo4eva):

the correct way this goes is 3x+6=5x-4

OpenStudy (jojo4eva):

you subtract 3x from both sides and you get 6=2x-4. you add 4 to both sides and get 10=2x divide both sides by 2 and you get 5

OpenStudy (jojo4eva):

@hmichelle your answer is indeed incorrect. the answer is 5

OpenStudy (anonymous):

But if there is a negative in the equation wouldn't it be -5 anyway?

OpenStudy (anonymous):

yesssss @Telepoto101

OpenStudy (jojo4eva):

Ok then you can choose -5 but you will be wrong.

OpenStudy (anonymous):

its 5

OpenStudy (anonymous):

Oh it is 5!!

OpenStudy (anonymous):

I got all my problems correct!!

OpenStudy (jojo4eva):

thank you @kah.x the correct answer is @hmichelle you did the problem completely wrong

OpenStudy (anonymous):

Thanks!

OpenStudy (jojo4eva):

No problem @Telepoto101 but listen to me next time lol

OpenStudy (anonymous):

3x+6=5x-4 add 4 on both sides 3x+10=5x subtract 3x from both sides 10=2x divide by 2 5=x or x=5

OpenStudy (anonymous):

Lol! I will

OpenStudy (jojo4eva):

Alright now if you need any other help just tag me

OpenStudy (anonymous):

K

OpenStudy (jojo4eva):

@hmichelle you are fine we all make mistakes haha. I just wanted to make sure @Telepoto101 got the correct answer you are good tho